# Generalist Forecasting with Frozen Video Models via Latent Diffusion

### v1

@@ 7 Acknowledgments
We thank Ross Goroshin, Yana Hasson, Pauline Luc, Kevin Murphy, Mehdi S. M. Sajjadi, Sjoerd van Steenkiste, Andrew Zisserman, and Daniel Zoran for their insightful comments.
Appendix A Appendix
A.1 Ablations
Table 3: Diffusion vs Regression
Model Pixels Depth Point Tracks Box Tracks
4DS-e Regression 18.62 18.62 56.23 0.193 0.193 784.3 0.59 0.59 0.00070 0.58 0.58 2.26
4DS-e 17.94 19.66 32.26 0.181 0.096 599.43 0.58 0.61 0.00068 0.56 0.66 1.87
WALT Regression 19.38 19.38 15.29 0.200 0.200 226.45 0.61 0.61 0.00140 0.54 0.54 2.44
WALT 500M 18.19 19.53 8.60 0.230 0.130 228.92 0.64 0.68 0.00134 0.50 0.58 2.47
Future events are not certain; they often require modeling of stochasticity. To account for this stochasticity, we propose metrics such as best of N and Frechet Distance. A deterministic regressor may be optimal in predicting the mean outcome, but it fails to account for the variance in possible outcomes. We believe these distributional metrics are a better measure of performance than simple mean prediction.
To demonstrate this, we perform an ablation of our proposed diffusion forecasting module and compare it to regression-based forecasting in Table 4. We find that while regression optimizes over the traditional mean-based metric, this does not account for the inherent stochasticity of forecasting. When considering metrics such as Frechet Distance or Best-of-N that take this stochasticity into account, we find that diffusion models generally outperform the regression baselines on in most cases.
A.2 Results Appendix
@@
Table 5: Model Performance Metrics
Model Pixel PSNR $\uparrow$ Depth Absolute Relative Error $\downarrow$ Jaccard Distance $\uparrow$ IoU $\uparrow$
Mean Std Max Perception Mean Std Min Perception Mean Std Max Perception Mean Std Max Perception
DINOv2 15.28 0.53 16.14 19.25 0.15 0.050 0.077 0.096 0.28 0.066 0.39 0.49 0.37 0.057 0.46 0.52
SigLIP 13.47 0.26 13.90 16.36 0.21 0.043 0.143 0.143 0.37 0.037 0.42 0.31 0.14 0.070 0.24 0.20
VideoPrism 16.54 0.53 17.39 21.85 0.16 0.049 0.081 0.102 0.54 0.049 0.61 0.74 0.44 0.045 0.51 0.63
VJEPA 17.12 0.66 18.11 20.95 0.20 0.052 0.119 0.107 0.59 0.023 0.63 0.79 0.49 0.054 0.58 0.74
VideoMAE 17.48 0.96 18.99 23.95 0.18 0.060 0.092 0.108 0.60 0.026 0.63 0.79 0.56 0.038 0.62 0.74
VideoMAEv2 16.55 0.87 17.92 21.82 0.20 0.069 0.093 0.113 0.56 0.040 0.61 0.73 0.56 0.053 0.64 0.72
4DS-h 17.78 1.17 19.64 24.62 0.21 0.070 0.110 0.120 0.40 0.110 0.58 0.82 0.43 0.130 0.63 0.74
4DS-e 17.94 1.08 19.66 26.03 0.18 0.056 0.096 0.091 0.58 0.029 0.61 0.83 0.56 0.071 0.66 0.77
WALT 500M 18.19 0.85 19.53 24.47 0.23 0.070 0.130 0.164 0.64 0.030 0.68 0.76 0.50 0.050 0.58 0.68
N-WALT 500M 20.59 0.11 20.77 24.47 0.24 0.031 0.186 0.164 0.50 0.020 0.53 0.76 0.41 0.039 0.47 0.68
@@
Appendix B Broader impacts
@@
Appendix C Timescales in Dataset Evaluation
We utilize 3 datasets for evaluation in the paper. They are ScanNet Dai et al. (2017), Perception Test Pătrăucean et al. (2023) and the Waymo Open Box dataset Sun et al. (2020). For all of them, we use the general setup in Carreira et al. (2024). The number of frames forecasted is always 12, but the effective sampling rate differs on each dataset. The range of forecasting thus varies from roughly half a second up to 3 seconds. On the Waymo dataset, the videos are sampled at 5 frames per second. On ScanNet, sampling is approximately 25 frames per second. Perception Test videos are sampled at a stride of 4.

### v2

@@ 7 Acknowledgments
We thank Ross Goroshin, Yana Hasson, Pauline Luc, Kevin Murphy, Mehdi S. M. Sajjadi, Sjoerd van Steenkiste, Andrew Zisserman, Neerja Thakkar, and Daniel Zoran for their insightful comments.
Appendix 0.A Appendix
Figure 4: Forecasting performance correlates with perceptual ability. We evaluate forecasting on pixels, point tracks, bounding boxes, and depth using 10 samples per example and report the normalized max (min for lower-is-better metrics) performance per task. We compare with models’ performance on the perception-style variant of each task on observed frames. Data points correspond to model types. In general, we find a linear correlation between forecasting and perception performance on these short-horizon tasks. However, for the best performing models this relationship is somewhat more complex.
0.A.1 Details on the Fréchet Distance Computation
Let $S_{g}=\{g_{1},g_{2},\dots,g_{n}\}$ be the set of $n$ ground truth trajectories and $S_{p}=\{p_{1},p_{2},\dots,p_{m}\}$ be the set of $m$ predicted trajectories. We first represent each trajectory as a vector in a $d$-dimensional space, where the dimensionality depends on the specific forecasting task:
  1. 1.
    Point tracks. Each trajectory consists of the 2D coordinates of a point over 12 future frames. By concatenating these coordinates, we form a vector $v_{i}\in\mathbb{R}^{24}$.
  2. 2.
    Box tracks. Each trajectory represents the 4 coordinates of an axis-aligned 2D bounding box over 12 future frames. This results in a vector $v_{i}\in\mathbb{R}^{48}$.
  3. 3.
    Depth and pixel predictions. Each trajectory is a sequence of 12 dense prediction frames. We downsample each frame to $14\times 14$ and concatenate them, yielding a vector $v_{i}\in\mathbb{R}^{2352}$.
For a given task, all ground truth and predicted trajectories are thus represented as vectors in the same space $\mathbb{R}^{d}$.
After collecting the ground truth and predicted trajectories, we model their distributions by fitting a multivariate Gaussian to each set. Specifically, we estimate the parameters of the ground truth distribution $\mathcal{N}_{g}(\mu_{g},\Sigma_{g})$ using the sample mean and unbiased sample covariance from the set $S_{g}$:
$\mu_{g}=\frac{1}{n}\sum_{i=1}^{n}g_{i}$
$\Sigma_{g}=\frac{1}{n-1}\sum_{i=1}^{n}(g_{i}-\mu_{g})(g_{i}-\mu_{g})^{T}$
The parameters for the predicted distribution, $\mu_{p}$ and $\Sigma_{p}$ of $\mathcal{N}_{p}(\mu_{p},\Sigma_{p})$, are computed similarly from the set of predicted trajectories $S_{p}$.
The Fréchet Distance (FD) is then defined as the squared Wasserstein-2 distance between these two estimated Gaussian distributions, and is calculated using the formula from [dowson1982frechet]:
[…]
where $||\cdot||_{2}^{2}$ denotes the squared Euclidean norm and $\text{Tr}(\cdot)$ is the trace of a matrix. The term $(\Sigma_{g}\Sigma_{p})^{1/2}$ represents the matrix square root of the product of the two covariance matrices. This metric provides a single scalar value that quantifies the dissimilarity between the distribution of predicted trajectories and that of the ground truth.
0.A.2 Pretraining Specs for Backbone Models
Table 3 lists the pretraining specs for all the vision models under consideration.
Model Type Variant Resolution Params Training Time
DINOv2 [oquab2023dinov2] I DINOV2_L/14 224 303M 637.5k steps
SigLIP [zhai2023siglip] I/L PaLI3-SigLIP-G-opt/14 224 2B 107k steps
VideoPrism [zhao2024videoprism] V/M/L v_giant 288 1.1B 500k steps
VJEPA [bardes2024vjepa] V/M L 224 300M 90k steps
VideoMAE [tong2022videomae] V/M H 224 600M 1600 epochs
VideoMAEv2 [wang2023videomaev2] V/M g 224 1B 1200 epochs
4DS-h [carreira2024scaling] V/M h 224 639M 488,282 steps
4DS-e [carreira2024scaling] V/M e 224 4B 488,282 steps
WALT [gupta2024walt] V/S/L 419M 128 419M 528k steps
0.A.3 Results Appendix
0.A.3.1 Forecasting Per-example Metrics
@@
Model Pixel PSNR $\uparrow$ Depth Abs. Rel. Error $\downarrow$
Mean Std Max Perception Mean Std Min Perception
DINOv2 16.38 0.857 17.78 19.97 0.146 0.038 0.088 0.089
SigLIP 13.64 0.39 14.29 16.48 0.233 0.036 0.175 0.144
VideoPrism 17.63 1.03 19.32 23.33 0.155 0.041 0.093 0.098
VJEPA 19.51 1.14 21.26 24.41 0.2 0.069 0.098 0.132
VideoMAE 20.23 1.1834 22.13 28.54 0.189 0.064 0.091 0.104
VideoMAEv2 18.59 1.41 20.81 26.44 0.206 0.071 0.1 0.107
4DS-h 20.54 1.40 22.69 31.24 0.224 0.077 0.107 0.1156
4DS-e 19.89 1.42 22.03 32.26 0.1937 0.065 0.096 0.086
WALT 500M 20.4 1.37 22.55 28.95 0.249 0.072 0.138 0.158
N-WALT 500M 22.48 0.19 22.78 28.95 0.19 0.026 0.145 0.158
Model Jaccard Distance $\uparrow$ IoU $\uparrow$
Mean Std Max Perception Mean Std Max Perception
DINOv2 0.28 0.066 0.39 0.49 0.37 0.057 0.46 0.52
SigLIP 0.37 0.037 0.42 0.31 0.14 0.070 0.24 0.20
VideoPrism 0.54 0.049 0.61 0.74 0.44 0.045 0.51 0.63
VJEPA 0.59 0.023 0.63 0.79 0.49 0.054 0.58 0.74
VideoMAE 0.60 0.026 0.63 0.79 0.56 0.038 0.62 0.74
VideoMAEv2 0.56 0.040 0.61 0.73 0.56 0.053 0.64 0.72
4DS-h 0.40 0.110 0.58 0.82 0.43 0.130 0.63 0.74
4DS-e 0.58 0.029 0.61 0.83 0.56 0.071 0.66 0.77
WALT 500M 0.64 0.030 0.68 0.76 0.50 0.050 0.58 0.68
N-WALT 500M 0.57 0.012 0.59 0.76 0.46 0.027 0.51 0.68
@@
0.A.3.2 N-WALT Metrics Across Different Noise Levels
The WALT model was jointly trained for frame prediction to enable long video generation through autoregressive prediction. To mitigate the domain shift between training, where ground-truth long videos are available, and inference, where the model predicts subsequent frames based on its own generated output, WALT incorporates noise conditioning augmentation [ho2022cascaded] during training. This augmentation involves applying noise to the past-frames-based conditioning signal. Concretely, conditioning noise is added in accordance with a noise schedule, by sampling a noise level as $t_{n}\sim\mathcal{U}(0,t_{\text{max}})$, where $t_{\text{max}}=300$. We evaluated the performance of N-WALT by varying the conditioning noise level during inference across the range $0,t_{\text{max}}]$. The performance metrics are detailed in Tables 5 and 6, where bold values denote the best task-specific metrics presented in Tables. 4 and  2, respectively.
Noise level Pixel PSNR $\uparrow$ Depth Absolute Relative Error $\downarrow$ Jaccard Distance $\uparrow$ IoU $\uparrow$
Mean Std Max Mean Std Min Mean Std Max Mean Std Max
0 21.46 0.310 21.95 0.237 0.032 0.186 0.500 0.020 0.533 0.407 0.039 0.471
10 22.48 0.188 22.78 0.190 0.018 0.160 0.564 0.010 0.580 0.462 0.027 0.507
20 22.38 0.199 22.70 0.189 0.020 0.156 0.566 0.010 0.582 0.455 0.030 0.504
30 22.33 0.214 22.67 0.189 0.022 0.153 0.565 0.012 0.584 0.453 0.032 0.506
40 22.37 0.230 22.74 0.187 0.024 0.149 0.571 0.012 0.590 0.454 0.034 0.510
50 22.39 0.249 22.79 0.186 0.026 0.145 0.570 0.014 0.593 0.457 0.037 0.517
100 22.19 0.331 22.73 0.188 0.036 0.132 0.562 0.020 0.593 0.447 0.048 0.524
150 21.88 0.412 22.54 0.193 0.044 0.124 0.570 0.025 0.610 0.439 0.056 0.531
200 21.59 0.488 22.37 0.194 0.053 0.113 0.553 0.031 0.602 0.433 0.062 0.534
250 21.30 0.561 22.19 0.200 0.059 0.109 0.553 0.037 0.610 0.428 0.066 0.535
300 20.94 0.635 21.95 0.205 0.066 0.104 0.550 0.041 0.611 0.425 0.070 0.539
Table 5: Per-example metrics by varying the conditioning noise level.
Noise level Pixels (ScanNet) Depth (ScanNet) Points (Perc. Test) Boxes (Waymo)
[…]
0 8.09 4.69 250.98 1.15 1.563 0.0378 2.72 0.0427
10 6.55 5.38 223.99 3.73 1.544 0.0379 3.23 0.0548
20 6.80 5.31 222.91 4.04 1.482 0.0378 2.89 0.0472
30 6.86 5.28 221.39 4.08 1.577 0.0377 2.90 0.0465
40 6.76 5.34 222.35 3.92 1.387 0.0379 2.72 0.0432
50 6.64 5.40 217.80 4.27 1.480 0.0378 3.13 0.0489
100 6.74 5.28 219.83 4.07 1.448 0.0378 2.60 0.0412
150 7.14 5.12 229.02 3.59 1.525 0.0378 2.66 0.0416
200 7.43 4.97 220.84 4.10 1.540 0.0378 2.78 0.0465
250 7.70 4.81 228.10 3.37 1.467 0.0378 2.89 0.0431
300 8.09 4.61 226.96 3.35 1.646 0.0369 3.04 0.0506
Table 6: Distribution-level metrics by varying the conditioning noise level.
Appendix 0.B Broader impacts
@@
Appendix 0.C Timescales in Dataset Evaluation
We utilize 3 datasets for evaluation in the paper. They are ScanNet [dai2017scannet], Perception Test [patraucean2023perception] and the Waymo Open Box dataset [Sun_2020_CVPR]. For all of them, we use the general setup in [carreira2024scaling]. The number of frames forecast is always 12, but the effective sampling rate differs on each dataset. The range of forecasting thus varies from roughly half a second up to 3 seconds. On the Waymo dataset, the videos are sampled at 5 frames per second. On ScanNet, sampling is approximately 25 frames per second. Perception Test videos are sampled at a stride of 4.
Model Pixels (ScanNet) Depth (ScanNet) Points (Perc. Test) Boxes (Waymo)
VideoMAEv1 6.35** 5.98** 2.34*** 9.13†
VideoMAEv2 23.85** 23.35** 6.33‡ 6.17‡
DinoV2 12.83** 12.48** 16.62† 16.55†
VideoPrism 14.90** 13.85** 3.98‡ 3.86‡
SigLip 21.74** 21.56** 18.62† 8.01‡
VJEPA 31.08** 31.05** 9.19† 9.13†
4DS-H 23.25** 22.38** 5.05§ 4.04§
4DS-e 45.41** 45.46** 13.14† 13.11†
Walt 45.75* 45.77* 13.71† 13.76†
Native Walt 9.78** 8.99** 9.63† 9.00**
Table 7: Experiment training time in hours. Both readout head and diffusion model were trained in parallel, but not end-to-end (stop gradient). Models were trained on TPUs. * represents 16 v5p chips, ** represents 4 v5p, *** 6 v5p, †4 v6e, ‡16 v6e, and §32 v5p.
Model Pixels Depth
Mean $\uparrow$ Best $\uparrow$ FD $\downarrow$ Mean $\downarrow$ Best $\downarrow$ FD
4DS-e 5 Samples 20.09 21.76 31.74 0.193 0.119 538.54
4DS-e 10 Samples 19.89 22.03 30.95 0.194 0.096 533.00
4DS-e 15 Samples 20.02 22.36 29.85 0.194 0.087 538.24
Table 8: Ablation of performance at different number of samples per example.
Model Pixels Depth
Perception Mean $\uparrow$ Best $\uparrow$ FD $\downarrow$ Perception Mean $\downarrow$ Best $\downarrow$ FD
4DS-e 512 29.99 20.40 22.39 29.74 0.087 0.187 0.096 527.72
4DS-e 1024 32.26 19.89 22.03 30.95 0.086 0.194 0.096 533.00
4DS-e 2048 31.62 20.50 22.58 30.46 0.088 0.192 0.093 530.64
Table 9: Ablation of performance at different capacities of the transformer-based readout head. 1024 is the default number used in the paper.